# SVBench: A Benchmark with Temporal Multi-Turn Dialogues for Streaming Video Understanding

### v1

@@ Abstract
Despite the significant advancements of Large Vision-Language Models (LVLMs) on established benchmarks, there remains a notable gap in suitable evaluation regarding their applicability in the emerging domain of long-context streaming video understanding. Current benchmarks for video understanding typically emphasize isolated single-instance text inputs and fail to evaluate the capacity to sustain temporal reasoning throughout the entire duration of video streams. To address these limitations, we introduce SVBench, a pioneering benchmark with temporal multi-turn question-answering chains specifically designed to thoroughly assess the capabilities of streaming video understanding of current LVLMs. We design a semi-automated annotation pipeline to obtain 49,979 Question-Answer (QA) pairs of 1,353 streaming videos, which includes generating QA chains that represent a series of consecutive multi-turn dialogues over video segments and constructing temporal linkages between successive QA chains. Our experimental results, obtained from 14 models in dialogue and streaming evaluations, reveal that while the closed-source GPT-4o outperforms others, most open-source LVLMs struggle with long-context streaming video understanding. We also construct a StreamingChat model, which significantly outperforms open-source LVLMs on our SVBench and achieves comparable performance on diverse vision-language benchmarks. We expect SVBench to advance the research of streaming video understanding by providing a comprehensive and in-depth analysis of current LVLMs. Our benchmark and model can be accessed at https://yzy-bupt.github.io/SVBench.
@@ 1 Introduction
In recent years, the rapid advancements in Large Language Models (LLMs) Touvron et al. (2023); Achiam et al. (2023) and visual processors Radford et al. (2021); Dosovitskiy et al. (2020) have significantly enhanced the performance of Large Vision-Language Models (LVLMs) Zhu et al. (2023); Ataallah et al. (2024); Maaz et al. (2023). These powerful models have been instrumental in pushing the boundaries of artificial intelligence, showcasing exceptional progress in domains such as visual reasoning and dialogue, particularly in the field of video understanding.
@@
Furthermore, there is a growing interest in applying these advancements to the emerging field of streaming video understanding Qian et al. (2024); Chen et al. (2024a). In conventional video understanding tasks, models are given the entire video and can leverage past, current, and future content to comprehend a specific video segment. In real-world scenarios, such as live streaming and security monitoring, video content is continually streamed, necessitating that dialogues should update concurrently with the temporal flow of the video without knowing the future information. Recently, streaming videos have become increasingly popular with diverse sources, including online video platforms Zellers et al. (2022), live streaming services Gao et al. (2023), and wearable camera footage Grauman et al. (2022). Therefore, researchers have begun to study LVLMs with the ability to interpret and interact with streaming content Qian et al. (2024); Chen et al. (2024a). Given these research advancements, it becomes imperative to establish a comprehensive evaluation benchmark specifically tailored to assess the progress in streaming video understanding achieved by LVLMs.
@@
Existing video benchmarks primarily focus on single-turn text input and cannot measure the ability to conduct temporally sequential reasoning, which falls short of capturing the complexity of streaming video understanding. For instance, several question-answering benchmarks Yu et al. (2019); Zhang et al. (2023); Xiao et al. (2021) typically comprise disjointed QA pairs tied to individual video clips, ignoring the continuous and dynamic nature of video streams. However, in real-world scenarios, users usually ask multiple questions during the entire duration of a video stream, with potential relevance among questions and video clips. Such complex human-computer interaction necessitates mastering fundamental skills, including engaging in multi-turn dialogues and comprehending extensive contextual histories to foster coherent and contextually relevant conversations. Although some current benchmarks partially assess essential interaction abilities, they are limited by relying on either multiple static images Feng et al. (2022) or short clips Zhao et al. (2018) to simulate multi-turn dialogues, rather than using long streaming videos. Therefore, appropriate benchmarks are required to evaluate the ability of streaming video understanding. To tackle these problems, we propose the first benchmark for streaming video understanding, named Streaming Video understanding Benchmark (SVBench), which aims at comprehensively evaluating the temporal multi-turn dialogue capabilities of LVLMs for streaming videos.
@@
Third, we conduct extensive evaluations, in terms of dialogue and streaming performance, of various prevalent LVLMs on our SVBench. Our results provide the first overview insight into the streaming video understanding capability of existing LVLMs. Surprisingly, these state-of-the-art LVLMs are far from satisfactory, in terms of streaming video understanding. These results motivate us to develop a stronger LVLM, namely StreamingChat, which significantly improves the overall dialogue evaluation score by 9.41% and the dialogue evaluation score by 3.30% on our SVBench, compared to the top-performing open-source LVLMs, while achieving comparable performance on conventional image and video benchmarks. Our benchmark and model will be publicly available, in order to catalyze the progress in streaming video understanding.
@@
EgoSchema Mangalam et al. (2024) 5,063 1.00 Auto&Manual
ActivityNet-QA Yu et al. (2019) 800 1.00 Manual
MVBench Li et al. (2024) 4,000 1.10 Auto
How2QA Li et al. (2020) 44,007 2.00 Manual
Perception Test Patraucean et al. (2024) 44,000 3.79 Auto&Manual
Social-IQ Zadeh et al. (2019) 7,500 6.00 Auto&Manual
MSVD-QA Xu et al. (2017) 13,157 6.68 Auto
TVQA Lei et al. (2018) 152,545 7.00 Manual
NExT-QA Xiao et al. (2021) 52,044 9.57 Manual
MovieChat Song et al. (2024) 13,000 13.00 Manual
LVBench Wang et al. (2024b) 1,549 15.04 Manual
TGIF-QA Jang et al. (2017) 165,165 17.25 Auto&Manual
MSRVTT-QA Xu et al. (2017) 72,821 24.35 Auto
@@ 2 Related Work
2.1 Large Vision-Language Models for Video
Recent advancements in Large Language Models (LLMs) have paved the way for a significant research focus on Large Vision-Language Models (LVLMs) aimed at improving multimodal understanding, such as multimedia retrieval Yang et al. (2024b; a), particularly in video content. Currently, in addition to the popular closed-source LVLMs such as GPT-4V Yang et al. (2023), GPT-4o Achiam et al. (2023), and Gemini 1.5 Pro Reid et al. (2024), an increasing number of open-source LVLMs, including Video-ChatGPT Maaz et al. (2023), VideoLLaMA2 Cheng et al. (2024), and VILA Lin et al. (2024), also have demonstrated the impressive capability in video understanding tasks. Advanced human-computer interaction in everyday life requires the ability to engage in multi-turn dialogues and understand extensive contextual histories to maintain coherent and contextually appropriate conversations Qian et al. (2024); Chen et al. (2024a). However, these LVLMs are still not fully adept at handling the intricacies of streaming videos and do not completely grasp the complexities of real-world contexts. To rigorously evaluate the capabilities of these models, we propose SVBench to measure the performance of LVLMs in video-related tasks that imitate the complexity of real-world interactions.
2.2 Video Understanding Benchmarks
In recent years, the exponential growth of video data has elevated video understanding to a crucial area within computer vision. To rigorously assess the performance of LVLMs, researchers have introduced a range of standardized benchmarks. These benchmarks provide not only comparative evaluation criteria for models but also catalyze advancements in video understanding models. The recent benchmarks, such as TGIF-QA Jang et al. (2017), MSVD-QA Xu et al. (2017), and MVBench Li et al. (2024), primarily comprise relatively brief videos capturing single events, thereby overlooking the temporal dependencies inherent in longer videos. To address the intricacies of long video comprehension, benchmarks using longer videos like movies have been developed. For instance, LLaMA-Vid Li et al. (2023), based on MovieNet Huang et al. (2020), has developed a movie QA dataset to identify character relationships. Similarly, MovieChat Su et al. (2020) employs a diverse set of videos and avoids specific character names or plot details within its questions. However, these benchmarks often fall short in addressing the intricate challenges posed by streaming videos, which encompass extended temporal contexts and dynamic scene variations. Therefore, we establish SVBench, a novel and comprehensive benchmark that aims to bridge this gap by offering an elaborate evaluation framework for long-context streaming video understanding.
@@ 3.1 Data Filtering and Scene Splitting
[…]
We propose a semi-automated annotation pipeline for streaming videos, as shown in Figure 2, including a multi-stage LLM-assisted generation process with several rounds of manual annotation. See Appendix G for detailed prompts. The annotation takes about 3 months and involves over 30 professional annotators.
@@ 3.1.1 Constructing QA Chains for Video Dialogues
[…]
@@ 5 StreamingChat
Model Architecture. Built upon InternVL2 Chen et al. (2024d), we develop a streaming LVLM baseline named StreamingChat. It comprises a vision encoder (InternViT Chen et al. (2023)), an MLP projector, and an LLM (InternLM2 Cai et al. (2024)), as illustrated in Figure 4. For the vision encoder, we employ the InternViT model (pre-trained on a combination of image captioning and OCR-specific datasets) to extract video frame embeddings at 1 FPS. To enhance the efficiency of learning streaming video understanding capabilities, we fine-tune the model using a static resolution strategy, which allows the model to handle several minutes of video and context within a 32k context window. The extracted frame embeddings are then fed into an MLP projector to generate frame tokens, following the approach used in LLaVA-1.5 Liu et al. (2024a). These frame tokens are interleaved with language tokens and input into the LLM, InternLM2. Additionally, we incorporate LoRA Hu et al. (2021) in every linear layer of the LLM to facilitate efficient tuning.
@@ 6.2 Evaluation Metrics
  • METEOR Banerjee & Lavie (2005) evaluates the precision, recall, and alignment of words and phrases between the references and the ground truth by considering synonymy and stemming.
@@
Dialogue Evaluation Framework. To evaluate the capabilities in both dialogue and streaming evaluations, it is crucial to adopt a multidimensional framework that assesses the quality of multi-turn dialogues. We propose an LLM-based evaluation framework encompassing several key aspects that contribute to the holistic assessment of LVLMs. See Appendix G for detailed prompts.
@@
MovieChat Song et al. (2024) 21.91 21.32 18.82 17.78 24.58 19.23 16.77 16.59 15.41
Video-ChatGPT Maaz et al. (2023) 35.85 40.12 31.18 29.37 39.38 29.95 25.77 26.97 22.79
TimeChat Ren et al. (2024) 33.26 43.62 29.86 30.58 39.97 28.49 27.08 28.10 24.22
VideoLLaMA2 Cheng et al. (2024) 36.76 47.38 30.56 33.08 39.85 37.09 29.67 31.75 26.03
ShareGPT4Video Chen et al. (2024b) 39.92 47.54 33.61 33.56 43.69 35.97 28.98 31.25 26.43
Video-LLaVA Lin et al. (2023) 38.99 46.42 35.86 34.27 46.09 35.65 31.21 31.20 26.67
InternVL2 Chen et al. (2024d) 37.10 51.19 38.25 36.79 36.02 41.66 32.22 33.24 29.07
VILA Lin et al. (2024) 42.13 55.48 40.60 39.48 45.93 43.29 35.87 34.22 31.21
InternLM-XComposer2.5 Zhang et al. (2024b) 47.64 57.71 40.49 41.72 50.00 47.20 36.99 39.53 33.90
MiniCPM-V 2.6 Yao et al. (2024) 43.25 50.45 39.93 39.96 51.61 47.40 37.31 38.67 33.43
@@
Gemini 1.5 Pro Reid et al. (2024) 41.29 43.32 42.24 39.10 47.20 50.80 39.19 34.78 35.98
GPT-4V Yang et al. (2023) 56.84 61.57 49.10 51.31 58.57 55.76 47.47 47.44 42.97
GPT-4o Achiam et al. (2023) 57.95 59.47 52.29 49.97 56.63 58.65 49.02 47.47 44.58
@@ 6.3 Overall Performance
Additionally, we conduct performance comparisons before and after fine-tuning on 6 image and video understanding benchmarks, as illustrated in Figure 5. The results indicate that the fine-tuned StreamingChat shows substantial improvements over the original InternVL2 on SVBench. While there are slight decreases in performance on the image benchmark MMBench Liu et al. (2023) and the video benchmark MMBench-Video Fang et al. (2024), there are modest gains on the video benchmarks VideoMME Fu et al. (2024) and MVBench Li et al. (2024). These findings suggest that StreamingChat enhances streaming video understanding capabilities without compromising fundamental image and video comprehension skills.

### v2

@@ Abstract
Despite the significant advancements of Large Vision-Language Models (LVLMs) on established benchmarks, there remains a notable gap in suitable evaluation regarding their applicability in the emerging domain of long-context streaming video understanding. Current benchmarks for video understanding typically emphasize isolated single-instance text inputs and fail to evaluate the capacity to sustain temporal reasoning throughout the entire duration of video streams. To address these limitations, we introduce SVBench, a pioneering benchmark with temporal multi-turn question-answering chains specifically designed to thoroughly assess the capabilities of streaming video understanding of current LVLMs. We design a semi-automated annotation pipeline to obtain 49,979 Question-Answer (QA) pairs of 1,353 streaming videos, which includes generating QA chains that represent a series of consecutive multi-turn dialogues over video segments and constructing temporal linkages between successive QA chains. Our experimental results, obtained from 14 models in dialogue and streaming evaluations, reveal that while the closed-source GPT-4o outperforms others, most open-source LVLMs struggle with long-context streaming video understanding. We also construct a StreamingChat model, which significantly outperforms open-source LVLMs on our SVBench and achieves comparable performance on diverse vision-language benchmarks. We expect SVBench to advance the research of streaming video understanding by providing a comprehensive and in-depth analysis of current LVLMs. Our benchmark and model can be accessed at https://github.com/sotayang/SVBench.
@@ 1 Introduction
In recent years, the rapid advancements in Large Language Models (LLMs) touvron2023llama; achiam2023gpt and visual processors radford2021learning; dosovitskiy2020image have significantly enhanced the performance of Large Vision-Language Models (LVLMs) zhu2023minigpt; ataallah2024minigpt4; maaz2023video. These powerful models have been instrumental in pushing the boundaries of artificial intelligence, showcasing exceptional progress in domains such as visual reasoning and dialogue, particularly in the field of video understanding.
@@
Furthermore, there is a growing interest in applying these advancements to the emerging field of streaming video understanding qian2024streaming; chen2024videollm. In conventional video understanding tasks, models are given the entire video and can leverage past, current, and future content to comprehend a specific video segment. In real-world scenarios, such as live streaming and security monitoring, video content is continually streamed, necessitating that dialogues should update concurrently with the temporal flow of the video without knowing the future information. Recently, streaming videos have become increasingly popular with diverse sources, including online video platforms zellers2022merlot, live streaming services gao2023livechat, and wearable camera footage grauman2022ego4d. Therefore, researchers have begun to study LVLMs with the ability to interpret and interact with streaming content qian2024streaming; chen2024videollm. Given these research advancements, it becomes imperative to establish a comprehensive evaluation benchmark specifically tailored to assess the progress in streaming video understanding achieved by LVLMs.
@@
Existing video benchmarks primarily focus on single-turn text input and cannot measure the ability to conduct temporally sequential reasoning, which falls short of capturing the complexity of streaming video understanding. For instance, several question-answering benchmarks yu2019activitynet; zhang2023movqa; xiao2021next typically comprise disjointed QA pairs tied to individual video clips, ignoring the continuous and dynamic nature of video streams. However, in real-world scenarios, users usually ask multiple questions during the entire duration of a video stream, with potential relevance among questions and video clips. Such complex human-computer interaction necessitates mastering fundamental skills, including engaging in multi-turn dialogues and comprehending extensive contextual histories to foster coherent and contextually relevant conversations. Although some current benchmarks partially assess essential interaction abilities, they are limited by relying on either multiple static images feng2022mmdialog or short clips zhao2018multi to simulate multi-turn dialogues, rather than using long streaming videos. Therefore, appropriate benchmarks are required to evaluate the ability of streaming video understanding. To tackle these problems, we propose the first benchmark for streaming video understanding, named Streaming Video understanding Benchmark (SVBench), which aims at comprehensively evaluating the temporal multi-turn dialogue capabilities of LVLMs for streaming videos.
@@
Third, we conduct extensive evaluations, in terms of dialogue and streaming performance, of various prevalent LVLMs on our SVBench. Our results provide the first overview insight into the streaming video understanding capability of existing LVLMs. Surprisingly, these state-of-the-art LVLMs are far from satisfactory, in terms of streaming video understanding. These results motivate us to develop a stronger LVLM, namely StreamingChat, which significantly improves the overall dialogue evaluation score by 9.41% and the streaming evaluation score by 3.30% on our SVBench, compared to the top-performing open-source LVLMs, while achieving comparable performance on conventional image and video benchmarks. Our benchmark and model will be publicly available, in order to catalyze the progress in streaming video understanding.
@@
EgoSchema mangalam2024egoschema 5,063 1.00 Auto&Manual
ActivityNet-QA yu2019activitynet 800 1.00 Manual
MVBench li2024mvbench 4,000 1.10 Auto
How2QA li2020hero 44,007 2.00 Manual
Perception Test patraucean2024perception 44,000 3.79 Auto&Manual
Social-IQ zadeh2019social 7,500 6.00 Auto&Manual
MSVD-QA xu2017video 13,157 6.68 Auto
TVQA lei2018tvqa 152,545 7.00 Manual
NExT-QA xiao2021next 52,044 9.57 Manual
MovieChat song2024moviechat 13,000 13.00 Manual
LVBench wang2024lvbench 1,549 15.04 Manual
TGIF-QA jang2017tgif 165,165 17.25 Auto&Manual
MSRVTT-QA xu2017video 72,821 24.35 Auto
@@ 2 Related Work
Large Vision-Language Models for Video
Recent advancements in Large Language Models (LLMs) have paved the way for a significant research focus on Large Vision-Language Models (LVLMs) aimed at improving multimodal understanding, such as multimedia retrieval yang2024ldre; yang2024semantic, particularly in video content. Currently, in addition to the popular closed-source LVLMs such as GPT-4V yang2023dawn, GPT-4o achiam2023gpt, and Gemini 1.5 Pro reid2024gemini, an increasing number of open-source LVLMs, including Video-ChatGPT maaz2023video, VideoLLaMA2 cheng2024videollama, and VILA lin2024vila, also have demonstrated the impressive capability in video understanding tasks. Advanced human-computer interaction in everyday life requires the ability to engage in multi-turn dialogues and understand extensive contextual histories to maintain coherent and contextually appropriate conversations qian2024streaming; chen2024videollm. However, these LVLMs are still not fully adept at handling the intricacies of streaming videos and do not completely grasp the complexities of real-world contexts. To rigorously evaluate the capabilities of these models, we propose SVBench to measure the performance of LVLMs in video-related tasks that imitate the complexity of real-world interactions.
Video Understanding Benchmarks
In recent years, the exponential growth of video data has elevated video understanding to a crucial area within computer vision. To rigorously assess the performance of LVLMs, researchers have introduced a range of standardized benchmarks. These benchmarks provide not only comparative evaluation criteria for models but also catalyze advancements in video understanding models. The recent benchmarks, such as TGIF-QA jang2017tgif, MSVD-QA xu2017video, and MVBench li2024mvbench, primarily comprise relatively brief videos capturing single events, thereby overlooking the temporal dependencies inherent in longer videos. To address the intricacies of long video comprehension, benchmarks using longer videos like movies have been developed. For instance, LLaMA-Vid li2023llama, based on MovieNet huang2020movienet, has developed a movie QA dataset to identify character relationships. Similarly, MovieChat su2020moviechats employs a diverse set of videos and avoids specific character names or plot details within its questions. However, these benchmarks often fall short in addressing the intricate challenges posed by streaming videos, which encompass extended temporal contexts and dynamic scene variations. Therefore, we establish SVBench, a novel and comprehensive benchmark that aims to bridge this gap by offering an elaborate evaluation framework for long-context streaming video understanding.
@@ 3.1 Data Filtering and Scene Splitting
We source 12,989 raw video data from a variety of publicly available datasets, including YT-Temporal-1B zellers2022merlot, YouCook2 zhou2018towards, ActivityNet caba2015activitynet, MovieChat song2024moviechat, Panda-70M chen2024panda, and Ego4D grauman2022ego4d. These datasets offer a wide spectrum of video content, ensuring a rich diversity that is essential for streaming video understanding. We then filter high-quality videos of adequate lengths, high aesthetic scores, and appropriate optical flow scores. Subsequently, we employ PySceneDetect [^1] https://github.com/Breakthrough/PySceneDetect to identify and enumerate scenes within filtered videos. These results are crucial in determining whether the video content exhibits adequate variation and complexity. Further filtering is conducted to retain only those videos containing 5 to 15 scenes, thereby excluding content that is either excessively monotonous or overly intricate. Moreover, only videos with an appropriate average scene duration are chosen, ensuring fluidity and rhythm. Finally, 1,353 videos are selected. Ultimately, we split each video into clips $V=\{s_{i}\ |\ 0\leq i<|V|\}$ based on timestamps, merging clips shorter than 2 seconds with their adjacent ones. Notably, to prevent disjointedness and ensure continuity between scenes, we add an extra 0.5 seconds to both the beginning and end of each clip, resulting in a one-second overlap between consecutive clips. More details are included in Appendix A.
3.2 Annotation Pipeline
We propose a semi-automated annotation pipeline for streaming videos, as shown in Figure 2, including a multi-stage LLM-assisted generation process with several rounds of manual annotation. See Appendix H for detailed prompts. The annotation takes about 3 months and involves over 30 professional annotators.
@@ 3.2.1 Constructing QA Chains for Video Dialogues
We propose creating a multi-turn question-answering dataset on videos to evaluate the ability of LVLMs to conduct multi-turn dialogues and comprehend video-related contextual information. This ability is essential for the coherence and relevance of conversations in various contexts. Initially, we define a series of multi-turn question-and-answer interactions on video clips as QA chains. For every video clip $s_{i}\in V$, we build a QA chain $C_{i}=\{\langle Q_{j}^{i},A_{j}^{i}\rangle\ |\ 0\leq j<|C_{i}|\}$, where $Q_{j}^{i}$ and $A_{j}^{i}$ represent the $j$-th question and answer generated on the $i$-th video clip. To achieve this, we harness the video understanding capabilities of existing LVLMs (e.g. GPT-4o):
@@ 5 StreamingChat
Model Architecture. Built upon InternVL2 chen2024far, we develop a streaming LVLM baseline named StreamingChat. It comprises a vision encoder (InternViT chen2023internvl), an MLP projector, and an LLM (InternLM2 cai2024internlm2), as illustrated in Figure 4. For the vision encoder, we employ the InternViT model (pre-trained on a combination of image captioning and OCR-specific datasets) to extract video frame embeddings at 1 FPS. To enhance the efficiency of learning streaming video understanding capabilities, we fine-tune the model using a static resolution strategy, which allows the model to handle several minutes of video and context within a 32k context window. The extracted frame embeddings are then fed into an MLP projector to generate frame tokens, following the approach used in LLaVA-1.5 liu2024improved. These frame tokens are interleaved with language tokens and input into the LLM, InternLM2. Additionally, we incorporate LoRA hu2021lora in every linear layer of the LLM to facilitate efficient tuning.
@@ 6.2 Evaluation Metrics
  • METEOR banerjee2005meteor evaluates the precision, recall, and alignment of words and phrases between the references and the ground truth by considering synonymy and stemming.
@@
Dialogue Evaluation Framework. To evaluate the capabilities in both dialogue and streaming evaluations, it is crucial to adopt a multidimensional framework that assesses the quality of multi-turn dialogues. We propose an LLM-based evaluation framework encompassing several key aspects that contribute to the holistic assessment of LVLMs. See Appendix H for detailed prompts.
@@
MovieChat song2024moviechat 21.91 21.32 18.82 17.78 24.58 19.23 16.77 16.59 15.41
Video-ChatGPT maaz2023video 35.85 40.12 31.18 29.37 39.38 29.95 25.77 26.97 22.79
TimeChat ren2024timechat 33.26 43.62 29.86 30.58 39.97 28.49 27.08 28.10 24.22
VideoLLaMA2 cheng2024videollama 36.76 47.38 30.56 33.08 39.85 37.09 29.67 31.75 26.03
ShareGPT4Video chen2024sharegpt4video 39.92 47.54 33.61 33.56 43.69 35.97 28.98 31.25 26.43
Video-LLaVA lin2023video 38.99 46.42 35.86 34.27 46.09 35.65 31.21 31.20 26.67
InternVL2 chen2024far 37.10 51.19 38.25 36.79 36.02 41.66 32.22 33.24 29.07
VILA lin2024vila 42.13 55.48 40.60 39.48 45.93 43.29 35.87 34.22 31.21
InternLM-XComposer2.5 zhang2024internlm 47.64 57.71 40.49 41.72 50.00 47.20 36.99 39.53 33.90
MiniCPM-V 2.6 yao2024minicpm 43.25 50.45 39.93 39.96 51.61 47.40 37.31 38.67 33.43
@@
Gemini 1.5 Pro reid2024gemini 41.29 43.32 42.24 39.10 47.20 50.80 39.19 34.78 35.98
GPT-4V yang2023dawn 56.84 61.57 49.10 51.31 58.57 55.76 47.47 47.44 42.97
GPT-4o achiam2023gpt 57.95 59.47 52.29 49.97 56.63 58.65 49.02 47.47 44.58
@@ 6.3 Overall Performance
Additionally, we conduct performance comparisons before and after fine-tuning on 6 image and video understanding benchmarks, as illustrated in Figure 5. The results indicate that the fine-tuned StreamingChat shows substantial improvements over the original InternVL2 on SVBench. While there are slight decreases in performance on the image benchmark MMBench liu2023mmbench and the video benchmark MMBench-Video fang2024mmbench, there are modest gains on the video benchmarks VideoMME fu2024video and MVBench li2024mvbench. These findings suggest that StreamingChat enhances streaming video understanding capabilities without compromising fundamental image and video comprehension skills.